The plasma density in the lower ionosphere has
been measured during satellite re-entry to be
about 1018 m-3 at 50km altitude, 1017 at 70 km,
and 1014 at 8km. What are the plasma frequencies
there?
1
Expert's answer
2020-07-03T10:03:37-0400

The expression for plasma density in the ionosphere depends on the concentration in the following way:


"f(n)=\frac{\omega(n)}{2\pi}=\frac{e}{2\pi}\sqrt{\frac{n}{m_e\epsilon_0}}."

This gives values of frequency of

89.8 MHz at 8 km,

8.98 GHz at 50 km,

2.84 GHz at 70 km.